OpenStudy (anonymous):

Choose the multiplicative inverse. 4 + 3i

OpenStudy (anonymous):

the snappy answer is \[\frac{1}{4+3i}\] but that is not in standard form.

OpenStudy (anonymous):

standard form is \[\frac{4}{25}-\frac{3}{25}i\]

OpenStudy (anonymous):

4-3i/25

OpenStudy (anonymous):

if you need steps let me know

OpenStudy (anonymous):

I'd like to know the steps.. I've never heard the term multipicative inverse before.

OpenStudy (anonymous):

well i did not asked the question too, but i'd like to see too :D

OpenStudy (anonymous):

multiplicative inverse of a complex number is that number which on multiplication with the original number would give you 1 for a complex number z multiplicative inverse is

OpenStudy (anonymous):

ok the trick is (unless you just want a formula) to multiply \[\frac{1}{4+3i}\times \frac{4-3i}{4-3i}\] the denominator is easy because \[(a+bi)(a-bi)=a^2+b^2\] and the numerator is easy also because \[1(a-bi)=a-bi\] so you get \[\frac{a-bi}{a^2+b^2}\] or \[\frac{a}{a^2+b^2}-\frac{b}{a^2+b^2}i\] in standard form

OpenStudy (anonymous):

so for example the multiplicative inverse of \[4+3i\] is \[\frac{4}{25}-\frac{3}{25}i\] and the multiplicative inverse of \[2-5i\] is \[\frac{2}{29}+\frac{5}{29}i\]

OpenStudy (anonymous):

thanks for remark :D